OpenStudy (anonymous):

If I was to put a nickel into a jar and then put in double the amount each year, how much money would I have after 10 years?

OpenStudy (anonymous):

would it be $51.15 ?

OpenStudy (anonymous):

this is a geometric series with a = 1 and r = 2 sum after 10 years = 1. 2^10 - 1 -------- 2 - 1 = 2^10 - 1 = 1023 nickels

OpenStudy (anonymous):

yay! So, im right!

OpenStudy (anonymous):

0.05*(1+2+2^2+2^3+...+2^9)=0.05*(2^10-1) dollars At least if nickel means 0.05 dollars...

OpenStudy (anonymous):

sorry but i dont know what a nickel is

OpenStudy (anonymous):

nickel = 5 cents?